OpenStudy (dmndlife24):

find zeros, x-int, y-int, and domain and the range of the function f(x) = 3x+25/20

OpenStudy (anonymous):

Find the zeros of the equation, by setting f(x)=0. 0=3x+25/20; 25/20=1.25 -25/20=3x, -1.25/3=x , x is approximately -0.41667.

OpenStudy (anonymous):

The domain of the function is (-infinity, -0.41667]U[-0.41667, +infinity) To explain this notation, it means the graph extends from - infinity to + infinity while it hits the x-axis at -0.41667. The U means union, which combination of the two intervals.